OpenStudy (anonymous):

Amanda drove 420 miles using 18 gallons of gas. At this rate, how many gallons of gas would she need to drive 357 miles?

OpenStudy (anonymous):

18/420=3/70 (gallons per mile) (3/70)(357)=15.3 gallons